# Stochastic Population Update Provably Needs An Archive in Evolutionary Multi-objective Optimization

### v1

@@ 3.1 SMS-EMOA with SPU
SMS-EMOA presented in Algorithm 1 is a popular MOEA, which employs non-dominated sorting and hypervolume indicator to update the population. It starts from an initial population of $\mu$ solutions (line 1). In each generation, it randomly selects a solution $\bm{x}$ from the current population (line 3) for reproduction. With probability $p_{c}$, it selects another solution $\bm{y}$ and applies one-point crossover on $\bm{x}$ and $\bm{y}$ to generate an offspring solution $\bm{x}^{\prime}$ (lines 4–7); otherwise, $\bm{x}^{\prime}$ is set as the copy of $\bm{x}$ (line 9). Note that one-point crossover selects a random point $i\in\{1,2,\ldots,n\}$ and exchanges the first $i$ bits of two parent solutions, which actually produces two new solutions, but the algorithm only picks the one that consists of the first part of the first parent solution and the second part of the second parent solution. Afterwards, bit-wise mutation flips each bit of $\bm{x}^{\prime}$ with probability $1/n$ to produce an offspring $\bm{x}^{\prime\prime}$ (line 11). Then, the worst solution in $P\cup\{\bm{x}^{\prime\prime}\}$, the union of the current population and offspring, is removed (line 12) using the Population Update of SMS-EMOA subroutine described in Algorithm 2. The subroutine first partitions the solution set $Q$ (where $Q=P\cup\{\bm{x}^{\prime\prime}\}$) into non-dominated sets $R_{1},R_{2},\ldots,R_{v}$ , where $R_{1}$ contains all non-dominated solutions in $Q$, and $R_{i}$ ($i\geq 2$) contains all non-dominated solutions in $Q\setminus\cup_{j=1}^{i-1}R_{j}$. A solution $\bm{z}\in R_{v}$ is then removed by minimizing $\Delta_{\bm{r}}(\bm{x},R_{v}):=HV_{\bm{r}}(R_{v})-HV_{\bm{r}}(R_{v}\setminus\{\bm{x}\})$, where $HV_{\bm{r}}(X)$ denotes the hypervolume of the solution set $X$ with respect to a reference point $\bm{r}\in\mathbb{R}^{m}$ ($\forall i,r_{i}\leq\min_{\bm{x}\in\mathcal{X}}f_{i}(\bm{x})$), i.e., the volume of the objective space between the reference point and the objective vectors of the solution set. A larger hypervolume indicates better approximation of the Pareto front in terms of convergence and diversity.
In Bian et al. (2024b), the SPU method is introduced and shown to be beneficial for the search of MOEAs. During population updates, SPU randomly selects a fixed proportion $p_{s}$ of the current population and the offspring solution(s) to directly survive into the next generation and the removed part is selected from the rest solutions. This implies that each solution, including the worst solution in the population, has at least a probability $p_{s}$ of surviving to the next generation. Specifically, SPU of SMS-EMOA as presented in Algorithm 3 is used to replace the original Population Update procedure in line 12 of Algorithm 1. Note that $p_{s}$ is set to $1/2$ in Bian et al. (2024b), while we consider a general $p_{s}$ here.
@@
1:  $P\leftarrow\mu$ solutions uniformly and randomly selected from $\{0,1\}^{\!n}$ with replacement;
@@
Output: $|Q|-1$ solutions from $Q$
@@
Output: $|Q|-1$ solutions from $Q$
1:  $Q^{\prime}\leftarrow\lfloor|Q|\cdot(1-p_{s})\rfloor$ solutions uniformly and randomly selected from $Q$ without replacement;
@@
The expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump has been proved to be $O(\mu n^{k}\cdot\min\{1,\mu/2^{k/4}\})$ Bian et al. (2024b), which is better than that, i.e., $O(\mu n^{k})$, of the original SMS-EMOA. Intuitively, by introducing randomness into the population update, the evolutionary search has a chance to go along inferior regions which are close to Pareto optimal regions, thereby making the search easier. Here, we re-prove a tighter upper bound on the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, as shown in Theorem 1. It is also more general, as it considers a survival probability $p_{s}\in[1/(\mu+1),1-1/o(\mu))$ and a crossover probability $p_{c}=O(1)$, rather than the specific probabilities $p_{s}=1/2$ and $p_{c}=0$ analyzed in Bian et al. (2024b). Note that the running time of EAs is often measured by the number of fitness evaluations, the most time-consuming step in the evolutionary process.
@@
For SMS-EMOA solving OneJumpZeroJump with $n-2k=\Omega(n)$, if using SPU with survival probability $p_{s}\in[1/(\mu+1),1-1/o(\mu))$, the crossover probability $p_{c}=O(1)$, and the population size $\mu\geq(n-2k+4)/(1-p_{s})$, then the expected running time for finding the whole Pareto front is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$.
The proof of Theorem 1 needs Lemma 2, which shows that given a proper value of $\mu$, an objective vector on the Pareto front will always be maintained once it has been found. The reason is that in Algorithm 3, the removed solution is selected from $\left\lfloor(\mu+1)\cdot(1-p_{s})\right\rfloor\geq n-2k+4$ solutions in $Q$. For each objective vector on the Pareto front, whose size is $n-2k+3$, only one solution has positive $\Delta$-value. This ensures that these solutions rank among the top $n-2k+3$ solutions, and thus will not be removed.
@@
For SMS-EMOA solving OneJumpZeroJump, if using SPU with survival probability $p_{s}\in[1/(\mu+1),1-1/o(\mu))$, and the population size $\mu\geq(n-2k+4)/(1-p_{s})$, then an objective vector $\bm{f}^{*}$ on the Pareto front will always be maintained once it has been found.
@@
Suppose the objective vector $(a,n+2k-a),a\in[2k\dots n]\cup\{k,n+k\}$, on the Pareto front is obtained by SMS-EMOA, i.e., there exists at least one solution in $Q$ (i.e., $P\cup\{\bm{x}^{\prime}\}$ in line 12 of Algorithm 1) corresponding to the objective vector $(a,n+2k-a)$. Because only one solution is removed in each generation by Algorithm 1, we only need to consider the case that exactly one solution (denoted as $\bm{x}^{*}$) corresponds to the objective vector $(a,n+2k-a)$. Since $\bm{x}^{*}$ cannot be dominated by any other solution, we have $\bm{x}^{*}\in R_{1}$ in the SPU of SMS-EMOA procedure. Moreover, $R_{1}$ cannot contain solutions whose number of 1-bits are in $[1..k-1]\cup[n-k+1..n-1]$, because these solutions must be dominated by $\bm{x}^{*}$. We also have $\Delta_{\bm{r}}(\bm{x}^{*},R_{1})=HV_{\bm{r}}(R_{1})-HV_{\bm{r}}(R_{1}\setminus\{\bm{x}^{*}\})>0$ because the region
@@
cannot be covered by any objective vector in $\bm{f}(\{0,1\}^{n})\setminus\{(a,n+2k-a)\}$. If at least two solutions in $R_{1}$ have the same objective vector, they must have a zero $\Delta$-value, because removing one of them will not decrease the hypervolume covered. Thus, for each objective vector $(b,n+2k-b),b\in[2k..n]\cup\{k,n+k\}$, at most one solution can have a $\Delta$-value larger than zero, implying that there exist at most $n-2k+3$ solutions in $R_{1}$ with $\Delta>0$. Note that the removed solution is chosen from $\left\lfloor(\mu+1)\cdot(1-p_{s})\right\rfloor\geq n-2k+4$ solutions in $Q$. Thus, $\bm{x}^{*}$ will not be removed because it is one of the best $n-2k+3$ solutions. ∎
@@
Since the objective vectors in $F_{I}^{*}$ will always be maintained by Lemma 2, we could use the conclusion of Theorem 1 in Bian et al. (2024b), which proves that the expected number of fitness evaluations of the first phase is $O(\mu(n\log n+k^{k}))$, where the term $O(\mu k^{k})$ is the expected number of fitness evaluations for finding one objective vector in $F_{I}^{*}$ when event $E$, i.e., any solution in the initial population has at most $(k-1)$ or at least $(n-k+1)$ 1-bits, happens. By Chernoff bound and $n-2k=\Omega(n)$, an initial solution has at most $(k-1)$ or at least $(n-k+1)$ 1-bits with probability $\exp(-\Omega(n))$, implying that event $E$ happens with probability $\exp(-\Omega(n))^{\mu}$. Thus, the term $O(\mu k^{k})$ can actually be omitted, implying that the expected number of fitness evaluations of the first phase is $O(\mu n\log n)$.
Next, we consider the second phase. By employing SPU, any solution (including dominated ones) can survive into the next generation with a probability of at least $p_{s}$. This means that the population can preserve some dominated solutions to gradually reach $1^{n}$. We assume a “jump” to be an event where a solution $\bm{x}$ with $|\bm{x}|_{1}\in[n-k..n-1]$ is selected, and a new dominated solution closer to $1^{n}$ is generated and preserved. Thus, $1^{n}$ can be reached more easily through multiple jumps across the gap of dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$). We refer to the dominated solutions along the multiple jumps as “stepping stones” and assume that $1^{n}$ can be reached through $M$ stepping stones. Let $p_{i}$ denote the probability that, after selecting the $i$-th stepping-stone solution as a parent, the next stepping-stone solution is successfully generated. Next, we consider consecutive $M+1$ jumps for generating $1^{n}$. Any failure during the intermediate jumps will result in restarting the process from the beginning. The first jump starts by selecting the solution $\bm{x}$ with $n-k$ 1-bits, generating the first stepping-stone solution, and preserving it, with a probability of at least $p_{0}p_{s}/\mu$. For the $i$-th jump ($11\leq i\leq M-$ 1), the $(i+1)$-th stepping-stone solution is generated from the previous one and preserved with a probability of at least $p_{i}p_{s}/\mu$. For the final jump, the Pareto optimal solution $1^{n}$ found must be preserved, and the jump occurs with a probability of $p_{M}/\mu$. Therefore, the consecutive $M+1$ jumps can happen with probability at least
@@
Then, we consider the probability $p_{i}$ of each successful jump. Let $k_{i}$ denote the distance of the jump from the $i$-th stepping-stone solution to the next one. That is, the $(i+1)$-th stepping-stone solution can be generated by flipping $k_{i}$ 0-bits of the $i$-th stepping-stone solution from the remaining $k-\sum_{j=0}^{i-1}k_{j}$ 0-bits while keeping the other bits unchanged. Thus, after selecting the $i$-th stepping-stone solution, the next stepping-stone solution can be generated if crossover is not performed (whose probability is $1-p_{c}$), and only $k_{i}$ 00-bits are flipped by bit-wise mutation (whose probability is $(1-1/n)^{n-k_{i}}\cdot\binom{k-\sum_{j=0}^{i-1}k_{j}}{k_{i}}/n^{k_{i}}$). We have
@@
where the crossover probability $p_{c}=O(1)$.
@@
Compared to the previous bound $O(\mu n^{k}\cdot\min\{1,\mu/2^{k/4}\})$ in Bian et al. (2024b), where SPU is used with $p_{s}=1/2$ and the crossover probability $p_{c}=0$, our bound $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ in Theorem 1 is tighter, where $C=e\mu/(p_{s}(1-p_{c}))$. When $p_{s}=1/2$ and $p_{c}=0$, our bound becomes $O(\mu n^{k}\cdot\min\{1,(e\ln(2e\mu)/k)^{k-1}\})$. Note that $(e\ln(2e\mu)/k)^{k-1}<\mu/2^{k/4}$, and an exponential improvement can be achieved when $k$ is large, e.g., $k=n/8$. The improvement is because in the second phase of finding the Pareto optimal solution $1^{n}$: 1) We consider $M+1$ “jumps” across the gap between dominated solutions, and account for all possible jump sizes to find $1^{n}$; 2) We select the optimal number of jumps, i.e., $M=\lceil(k-1)/\ln C-1\rceil$. Note that Bian et al. Bian et al. (2024b) only considered two fixed-size jumps.
Recently, Zheng and Doerr Zheng and Doerr (2024b) have extended SMS-EMOA to solve a many-objective problem, $m$OneJumpZero-Jump, and also derived the expected running time $O(\mu n^{k}\cdot\min\{1,\mu/2^{k}\})$ for solving OneJumpZeroJump. When $k>2e\ln(2e\mu)$, our bound $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$ brings an improvement ratio of $\Theta(\mu(k/(2e\ln(2e\mu)))^{k-1})$, which can also be exponential when $k$ is large, e.g., $k=n/8$. Compared to the analysis in Bian et al. (2024b), Zheng and Doerr Zheng and Doerr (2024b) considered all possible jump sizes. However, they also reduced the process to two “jumps”, whereas we consider the optimal number $M+1$ of “jumps” with $M=\lceil(k-1)/\ln C-1\rceil$. For $\log\mu\leq k<2e\ln(2e\mu)$, our bound doesn’t show superiority, which is because to minimize Eq. (6), we choose $M=\lceil(k-1)/\ln C-1\rceil$ to round up to the nearest integer, leading to an over-relaxation. If we instead set $M=1$, the same bound can be obtained.

### v2

@@ 3.1 SMS-EMOA with SPU
SMS-EMOA presented in Algorithm 1 is a popular MOEA, which employs non-dominated sorting and hypervolume indicator to update the population. It starts from an initial population of $\mu$ solutions (line 1). In each generation, it randomly selects a solution $\bm{x}$ from the current population (line 3) for reproduction. With probability $p_{c}$, it selects another solution $\bm{y}$ and applies one-point crossover on $\bm{x}$ and $\bm{y}$ to generate an offspring solution $\bm{x}^{\prime}$ (lines 4–7); otherwise, $\bm{x}^{\prime}$ is set as the copy of $\bm{x}$ (line 9). Note that one-point crossover selects a random point $i\in\{1,2,\ldots,n\}$ and exchanges the first $i$ bits of two parent solutions, which actually produces two new solutions, but the algorithm only picks the one that consists of the first part of the first parent solution and the second part of the second parent solution. Afterwards, bit-wise mutation flips each bit of $\bm{x}^{\prime}$ with probability $1/n$ to produce an offspring $\bm{x}^{\prime\prime}$ (line 11). Then, the worst solution in $P\cup\{\bm{x}^{\prime\prime}\}$, the union of the current population and offspring, is removed (line 12) using the Population Update of SMS-EMOA subroutine described in Algorithm 2. The subroutine first partitions the solution set $Q$ (where $Q=P\cup\{\bm{x}^{\prime\prime}\}$) into non-dominated sets $R_{1},R_{2},\ldots,R_{v}$ , where $R_{1}$ contains all non-dominated solutions in $Q$, and $R_{i}$ ($i\geq 2$) contains all non-dominated solutions in $Q\setminus\cup_{j=1}^{i-1}R_{j}$. A solution $\bm{z}\in R_{v}$ is then removed by minimizing $\Delta_{\bm{r}}(\bm{x},R_{v}):=HV_{\bm{r}}(R_{v})-HV_{\bm{r}}(R_{v}\setminus\{\bm{x}\})$, where $HV_{\bm{r}}(X)$ denotes the hypervolume of the solution set $X$ with respect to a reference point $\bm{r}\in\mathbb{R}^{m}$ ($\forall i,r_{i}\leq\min_{\bm{x}\in\mathcal{X}}f_{i}(\bm{x})$), i.e., the volume of the objective space between the reference point and the objective vectors of the solution set. A larger hypervolume indicates better approximation of the Pareto front in terms of convergence and diversity. For bi-objective problems, as defined in the original SMS-EMOA Beume et al. (2007), the algorithm omits the reference point and directly preserves the two boundary points, allowing the hypervolume to be calculated accordingly.
In Bian et al. (2024b), the SPU method is introduced and shown to be beneficial for the search of MOEAs. During population updates, SPU randomly selects a proportion $p_{s}$ of the current population and the offspring solution(s) to directly survive into the next generation and the removed part is selected from the rest solutions. This implies that each solution, including the worst solution in the population, has at least a probability $p_{s}$ of surviving to the next generation. Specifically, SPU of SMS-EMOA as presented in Algorithm 3 is used to replace the original Population Update procedure in line 12 of Algorithm 1. Note that $p_{s}$ is set to $1/2$ in Bian et al. (2024b), while we consider a general $p_{s}$ here.
@@
[…]
@@
Output: |Q|1|Q|-1 solutions from $Q$
@@
Output: |Q|1|Q|-1 solutions from $Q$
[…]
@@
The expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump has been proven to be $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ Bian et al. (2024b), which is better than that, i.e., $O(\mu n^{k})$, of the original SMS-EMOA. Intuitively, by by introducing randomness into the population update, the evolutionary search has a chance to go along inferior regions which are close to Pareto optimal regions, thereby making the search easier. Here, we re-prove a tighter upper bound on the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, as shown in Theorem 1. It is also more general, as it considers a survival probability $p_{s}\in[1/(\mu+1),1-o(1/\mu))$, rather than just $p_{s}=1/2$ as in Bian et al. (2024b). Note that the running time of EAs is often measured by the number of fitness evaluations.
@@
For SMS-EMOA solving OneJumpZeroJump with $n-2k=\Omega(n)$, if using SPU with survival probability $p_{s}\in[1/(\mu+1),1-o(1/\mu))$, the crossover probability $1-p_{c}=\Omega(1)$, and the population size $\mu\geq(n-2k+4)/(1-p_{s})$, then the expected running time for finding the whole Pareto front is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$.
The proof of Theorem 1 needs Lemma 2, which shows that given a proper value of $\mu$, an objective vector on the Pareto front will always be maintained once it has been found. The reason is that in Algorithm 3, the removed solution is selected from $\left\lfloor(\mu+1)\cdot(1-p_{s})\right\rfloor\geq n-2k+4$ solutions in $Q$. For each objective vector on the Pareto front, whose size is n2k+3n-2k+3, only one solution has positive $\Delta$-value. This ensures that these solutions rank among the top n2k+3n-2k+3 solutions, and thus will not be removed.
@@
For SMS-EMOA solving OneJumpZeroJump, if using SPU with survival probability $p_{s}\in[1/(\mu+1),1-o(1/\mu))$, and the population size $\mu\geq(n-2k+4)/(1-p_{s})$, then an objective vector $\bm{f}^{*}$ on the Pareto front will always be maintained once it has been found.
@@
Suppose the objective vector $(a,n+2k-a),a\in[2k\dots n]\cup\{k,n+k\}$, on the Pareto front is obtained by SMS-EMOA, i.e., there exists at least one solution in $Q$ (i.e., $P\cup\{\bm{x}^{\prime}\}$ in line 12 of Algorithm 1) corresponding to the objective vector (a,n+2ka)(a,n+2k-a). Because only one solution is removed in each generation by Algorithm 1, we only need to consider the case that exactly one solution (denoted as $\bm{x}^{*}$) corresponds to the objective vector (a,n+2ka)(a,n+2k-a). Since $\bm{x}^{*}$ cannot be dominated by any other solution, we have $\bm{x}^{*}\in R_{1}$ in the SPU of SMS-EMOA procedure. Moreover, $R_{1}$ cannot contain solutions whose number of 1-bits are in $1..k-1]\cup[n-k+1..n-1]$, because these solutions must be dominated by $\bm{x}^{*}$. We also have $\Delta_{\bm{r}}(\bm{x}^{*},R_{1})=HV_{\bm{r}}(R_{1})-HV_{\bm{r}}(R_{1}\setminus\{\bm{x}^{*}\})>0$ because the region
@@
cannot be covered by any objective vector in $\bm{f}(\{0,1\}^{n})\setminus\{(a,n+2k-a)\}$. If at least two solutions in $R_{1}$ have the same objective vector, they must have a zero $\Delta$-value, because removing one of them will not decrease the hypervolume covered. Thus, for each objective vector $(b,n+2k-b),b\in[2k..n]\cup\{k,n+k\}$, at most one solution can have a $\Delta$-value larger than zero, implying that there exist at most n2k+3n-2k+3 solutions in $R_{1}$ with $\Delta>0$. Note that the removed solution is chosen from $\left\lfloor(\mu+1)\cdot(1-p_{s})\right\rfloor\geq n-2k+4$ solutions in $Q$. Thus, $\bm{x}^{*}$ will not be removed because it is one of the best n2k+3n-2k+3 solutions. ∎
@@
Since the objective vectors in $F_{I}^{*}$ will always be maintained by Lemma 2, we could use the conclusion of Theorem 1 in Bian et al. (2024b), which proves that the expected number of fitness evaluations of the first phase is $O(\mu(n\log n+k^{k}))$, where the term $O(\mu k^{k})$ is the expected number of fitness evaluations for finding one objective vector in $F_{I}^{*}$ when event $E$, i.e., any solution in the initial population has at most (k1)(k-1) or at least (nk+1)(n-k+1) 1-bits, happens. By Chernoff bound and $n-2k=\Omega(n)$, an initial solution has at most (k1)(k-1) or at least (nk+1)(n-k+1) 1-bits with probability $\exp(-\Omega(n))$, implying that event $E$ happens with probability $\exp(-\Omega(n))^{\mu}$. Thus, the term $O(\mu k^{k})$ can actually be omitted, implying that the expected number of fitness evaluations of the first phase is $O(\mu n\log n)$.
Next, we consider the second phase. By employing SPU, any solution (including dominated ones) can survive into the next generation with a probability of at least $p_{s}$. This means that the population can preserve some dominated solutions to gradually reach $1^{n}$. We assume a “jump” to be an event where a solution $\bm{x}$ with $|\bm{x}|_{1}\in[n-k..n-1]$ is selected, and a new dominated solution closer to $1^{n}$ is generated and preserved. Thus, $1^{n}$ can be reached more easily through multiple jumps across the gap of dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$). We refer to the dominated solutions along the multiple jumps as “stepping stones” and assume that $1^{n}$ can be reached through $M$ stepping stones. Let $p_{i}$ denote the probability that, after selecting the $i$-th stepping-stone solution as a parent, the next stepping-stone solution is successfully generated. Next, we consider consecutive $M+1$ jumps for generating $1^{n}$. Any failure during the intermediate jumps will result in restarting the process from the beginning. The first jump starts by selecting the solution $\bm{x}$ with nkn-k 1-bits, generating the first stepping-stone solution, and preserving it, with a probability of at least $p_{0}p_{s}/\mu$. For the $i$-th jump ($1\leq i\leq M-1$), the $(i+1)$-th stepping-stone solution is generated from the previous one and preserved with a probability of at least $p_{i}p_{s}/\mu$. For the final jump, the Pareto optimal solution $1^{n}$ found must be preserved, and the jump occurs with a probability of $p_{M}/\mu$. Therefore, the consecutive $M+1$ jumps can happen with probability at least
@@
Then, we consider the probability $p_{i}$ of each successful jump. Let $k_{i}$ denote the distance of the jump from the $i$-th stepping-stone solution to the next one. That is, the $(i+1)$-th stepping-stone solution can be generated by flipping $k_{i}$ 0-bits of the $i$-th stepping-stone solution from the remaining $k-\sum_{j=0}^{i-1}k_{j}$ 0-bits while keeping the other bits unchanged. Thus, after selecting the $i$-th stepping-stone solution, the next stepping-stone solution can be generated if crossover is not performed (whose probability is $1-p_{c}$), and only $k_{i}$ 0-bits are flipped by bit-wise mutation (whose probability is $(1-1/n)^{n-k_{i}}\cdot\binom{k-\sum_{j=0}^{i-1}k_{j}}{k_{i}}/n^{k_{i}}$). We have
@@
where the crossover probability $1-p_{c}=\Omega(1)$.
@@
Under the same conditions as the previous results $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ in Bian et al. (2024b) and $O(\mu n^{k}\cdot\min\{1,\mu/2^{k}\})$ in Zheng and Doerr (2024b), where SPU is used with $p_{s}=1/2$ and crossover probability $p_{c}=0$, our bound in Theorem 1 becomes $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$. Since the bound in Zheng and Doerr (2024b) is tighter than that in Bian et al. (2024b), we focus on comparing our result only with Zheng and Doerr (2024b). When $k>2e\ln(2e\mu)$, our bound $O(\mu n^{k}\cdot(e\ln(2e\mu)/k)^{k-1})$ brings an improvement ratio of $\Theta(\mu(k/(2e\ln(2e\mu)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. For a very small range $\log\mu\leq k<2e\ln(2e\mu)$, our bound shows no advantage, which is because to minimize Eq. (6), we choose $M=\lceil(k-1)/\ln C-1\rceil$ to round up to the nearest integer, leading to an over-relaxation. If we instead set $M=1$, the same bound can be obtained. The reason of our better bound is in the second phase of finding the Pareto optimal solution $1^{n}$: 1) We consider $M+1$ “jumps” across the gap between dominated solutions, and account for all possible jump sizes to find $1^{n}$; 2) We select the optimal number of jumps, i.e., $M=\lceil(k-1)/\ln C-1\rceil$. Note that Bian et al. Bian et al. (2024b) considered only two fixed-size jumps, and although Zheng and Doerr Zheng and Doerr (2024b) accounted for all possible jump sizes, they also reduced the process to two jumps.